the sequence 6, 24, 54, 96,... comes from the surface area of cubes.
they want me to find the next number in the sequence but I don't know how.
I do know the formula for the surface area of a cube but I don't know what to plug in to get those numbers.

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

6, 24, 54, 96
surface area of cube = 6 *l^2 where l is the side of cube
6*1^2=6
6*2^2=24
6*3^2=54
6*4^2=96
6*5^2=150
6*6^2=216

Answer by stanbon(75887)

the sequence 6, 24, 54, 96,... comes from the surface area of cubes.
they want me to find the next number in the sequence but I don't know how.
------
If edge = 1unit, face = 1 unit^2, surface area = 6*1 = 6units^2
---
If edge = 2 unit, face = 4 unit^2, surface area = 6*4 = 24 units^2
---
If edge = 3 then face = 9 and surface area = 6*9 = 54 units^2
----
If edge = 4 then face = 16 and sa = 6*16 = 96 units^2
----
If edge = 5 then face = 25 a sa = 6*25 = 150 units^2
=============
Cheers,
Stan H.